### v1

# On Hamiltonian formulations of the Dirac system

Abstract
We extend a previously successful discussion of the constrained Schrödinger system through the Dirac–Bergmann algorithm to the case of the Dirac field. In order to follow the analogy, first we discuss the classical Dirac field as a spinorial variable, by introducing properly defined momenta and a suitably modified, factor ordered Poisson bracket. According to the Dirac–Bergmann algorithm two second class Hamiltonian constraints emerge, leading to a factor ordered Dirac bracket on the full phase space. This becomes the Poisson bracket on the reduced phase space in the canonical chart adapted to the shell. The Dirac equation is recovered both as consistency condition on the full phase space and as canonical equation on the reduced phase space. Alternatively, considering the Dirac field as odd Grassmann variable, we present the details of the Dirac–Bergmann algorithm (with either left and righ derivatives acting on Grassmann valued superfunctions and involving a different type of generalized Poisson and Dirac brackets). We propose a recipe for the canonical second quantization of all three versions of the generalized Dirac brackets, yielding the correct fundamental anticommutator.
I Introduction
Most of the physical theories of interest contain constraints, for which the transition from the Lagrangian to the Hamiltonian description requires the Dirac–Bergmann algorithm, developed initially by Dirac DiracConstr and Bergmann with his collaborators (for a historical account see Salisbury and for further details SM ; Sundermayer ; HT ). The full machinery of the formalism through the analysis of a finite-dimensional system has been recently presented in Ref. Brown .
While the corresponding discussions of the Maxwell electrodynamics and of general relativity Wald ; ADM are well known, the Schrödinger field as a constrained system was analyzed relatively late Schrod . There, the Schrödinger equation emerged both as a consistency condition in the full phase space and as canonical equation in the reduced phase space. By implementing the quantization scheme for systems with second class constraints, inconsistencies of previous treatments were eliminated.
In some respects the Dirac field is similar to the Schrödinger field. Most notably, its dynamical equation contains only first time derivatives (hence a constraint equation), moreover those appear only linearly, impending the Legendre transformation, and requiring the application of the Dirac–Bergmann algorithm. The Dirac field, being spinorial, is much more complicated, yielding anticommutators upon quantization. Hence, it is often considered an odd Grassmannian variable, in order to recover the same algebraic structure classically as in the quantum version of the theory.
In this paper, for the classical description of the Dirac field we attempt to follow as closely as possible the successful treatment of the Schrödinger field. We will show that the Dirac field classically can be characterized by a Hamiltonian structure through the algorithmic application of the Dirac–Bergmann procedure, similarly to the case of the Schrödinger field. Nevertheless, there is a price to pay due to the Dirac field being a spinor. In order the Dirac–Bergmann algorithm to work, we will generalize both the concepts of the Poisson and Dirac brackets such to observe a meaningful factor ordering. This has to be imposed during the canonical quantization as well.
The paper proceeds as follows. In Section 2 we introduce the Dirac spinor (also include a pedagogical discussion on its spin in the Appendix). Then, we review the variational procedures leading to the Dirac equation in a classical approach, simplified by the introduction of suitably defined ∂̸$\not{\partial}_{a}∂$ derivatives. The various Lagrangians for the Dirac field explored in the literature are analogous to the simplest Schiff and Hermitian Henley–Thirring Lagrangians for the Scrödinger field, discussed in Schrod . In contrast with the literature, in our discussion we define the canonical momenta as Dirac adjoint spinors, which will have multiple advantages.
Section 3 presents the Dirac–Bergmann algorithm applied to the Hermitian Lagrangian for the Dirac field. This leads to two primary Hamiltonian constraints, which are the Dirac adjoints of each other. We proceed with the definition of the canonical and primary Hamiltonians. Then we introduce a factor ordered Poisson bracket, in which Dirac adjoint spinors are always followed by spinors. Time evolution is defined through this Poisson bracket. The evolution of the constraints generates the Dirac equation and its Dirac adjoint as consistency conditions, holding weakly. The constraints are second class, which lead to the definition of the Dirac bracket, imposing again a factor ordering requirement, with the aim to have only Dirac adjoint spinor times spinor type inner products. Time evolution can be reexpressed in terms of the factor ordered Dirac bracket. The Dirac brackets of the canonical pairs contain $1/2$ factors not present in their Poisson brackets. This feature mimics closely the corresponding feature of the Schrödinger field Schrod , which turned advantageous in its canonical quantization. We identify the reduced phase space variables, proving that the Dirac bracket on the full phase space turns into the Poisson bracket on the reduced phase space. The Hamiltonian on the reduced phase space reproduces the Hermitian Hamiltonian evaluated on shell.
For comparing our approach with the interpretation of the classical Dirac field as an odd Grassmann variable, in Section 4 we also carry out the Dirac–Bergmann procedure with both left and righ derivatives acting on Grassmann valued superfunctions. These involve another type of generalized Poisson and Dirac brackets, which are symmetric, rather than antisymmetric, together with definitions of the momenta different from the literature. As far as we are aware of, this discussion was not presented in detail elsewhere.
In Section 5 we proceed with the canonical second quantization of all three versions of the constrained Dirac dynamics. The suitably defined quantization recipes transform all versions of Dirac brackets in a natural way into the fundamental anticommutator of the second quantized Dirac field.
[…]
II Standard variational procedures: a review
II.1 The Dirac equation
The Dirac spinor $\psi$ obeys the equation
$\left(i\frac{\hbar}{c}\gamma^{a}\partial_{a}-m\right)\psi=0~{},$ (1)
where $m$ is the mass and the Dirac matrices $\gamma^{a}$ are given through their anticommutators
$\left[\gamma^{a},\gamma^{b}\right]_{+}=-2\eta^{ab}\mathrm{I}_{4}\text{~{},}$ (2)
with $\eta_{ab}~{}$the Minkowski metric and $\mathrm{I}_{4}$ the unit matrix in 4 dimensions. In the Dirac representation they read as
$\gamma^{0}=\left(\begin{array}[]{cc}\mathrm{I}_{2}&0\\ 0&-\mathrm{I}_{2}\end{array}\right)~{},\quad\gamma^{\mu}=\left(\begin{array}[]{cc}0&\sigma^{\mu}\\ -\sigma^{\mu}&0\end{array}\right)~{},$ (3)
in the Weyl (chiral) representation as DWM
$\gamma_{W}^{0}=\left(\begin{array}[]{cc}0&\mathrm{I}_{2}\\ \mathrm{I}_{2}&0\end{array}\right)~{},\quad\gamma_{W}^{\mu}=\left(\begin{array}[]{cc}0&-\sigma^{\mu}\\ \sigma^{\mu}&0\end{array}\right)~{},$ (4)
while in the Majorana representation as DWM
$\gamma_{M}^{0}=\left(\begin{array}[]{cc}0&\sigma^{2}\\ \sigma^{2}&0\end{array}\right)~{},\quad\gamma_{M}^{1}=\left(\begin{array}[]{cc}i\sigma^{3}&0\\ 0&i\sigma^{3}\end{array}\right)~{},\quad\gamma_{M}^{2}=\left(\begin{array}[]{cc}0&-\sigma^{2}\\ \sigma^{2}&0\end{array}\right)~{},\quad\gamma_{M}^{3}=\left(\begin{array}[]{cc}-i\sigma^{1}&0\\ 0&-i\sigma^{1}\end{array}\right)~{}.$ (5)
(Note that the matrices $\gamma_{W}^{\mu}$ are sometimes defined with a global minus sign, see Ref. LancasterBlundell , which also equivalently defines $\gamma_{LB}^{1}=-\gamma_{M}^{3}$, $\gamma_{LB}^{2}=-\gamma_{M}^{2}$, and $\gamma_{LB}^{3}=\gamma_{M}^{1}$.)
In all representations the Dirac matrices are constructed in terms of the 2-dimensional unit matrix $\mathrm{I}_{2}$ and the Pauli matrices $\sigma^{\mu}$, which obey the commutation relations
$\left[\sigma^{\mu},\sigma^{\nu}\right]=2i\epsilon_{~{}~{}\rho}^{\mu\nu}\sigma^{\rho}~{},$ (6)
with $\epsilon_{\mu\nu\rho}$ the Levi-Civita symbol (its indices can freely raised and lowered due to the metric signature adopted). Explicitly the Pauli matrices are
$\sigma^{1}=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)~{},\quad\sigma^{2}=\left(\begin{array}[]{cc}0&-i\\ i&0\end{array}\right)~{},\quad\sigma^{3}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)~{},$ (7)
obeying $\left(\sigma^{\mu}\right)^{{\dagger}}=\sigma^{\mu}$ ( denotes the adjoint). Hence $\left(\gamma^{0}\right)^{{\dagger}}=\gamma^{0}$ and $\left(\gamma^{\mu}\right)^{{\dagger}}=-\gamma^{\mu}$ (identical relations hold for all representations). It is also useful to remember the relations
$\left(\gamma^{0}\right)^{{\dagger}}\gamma^{0}=\mathrm{I}_{4}~{},\quad\gamma^{0}\left(\gamma^{a}\right)^{{\dagger}}\gamma^{0}=\gamma^{a}.$ (8)
In the Appendix we present a simple proof that the Dirac spinor $\psi$ represents spin $1/2$ particles.
II.2 Lagrangian formalism with ∂̸$\not{\partial}_{a}∂$ derivatives
The action
[…]
characterizing the Dirac field depends on a pair of independent variables, the spinor $\psi$ and the Dirac adjoint spinor $\bar{\psi}=\psi^{{\dagger}}\gamma^{0}$. Here $\mathfrak{L}$ is the Lagrangian density, its derivative with respect to $\psi$ being a Dirac adjoint spinor, while with respect to $\bar{\psi}$ a spinor. The Dirac spinor is represented classically as a column matrix, while its Dirac adjoint spinor as a row matrix. Note that in order $\mathfrak{L}$ to be a Lorentz scalar, it has to contain products of Dirac adjoint spinors with spinors (in this order). Variation of the action leads to the Euler–Lagrange equations
$\frac{\partial\mathfrak{L}}{\partial\psi}-\partial_{a}\frac{\partial\mathfrak{L}}{\partial\partial_{a}\psi}=0~{},\qquad\frac{\partial\mathfrak{L}}{\partial\bar{\psi}}-\partial_{a}\frac{\partial\mathfrak{L}}{\partial\partial_{a}\bar{\psi}}=0~{}.$ (10)
As will be seen later, all Lagrangians leading to the Dirac equation contain only Feynman slash (Dirac slash) derivatives ∂̸$\not{\partial}=\gamma^{a}\partial_{a}∂$ (with summation over repeated indices understood). We further introduce the set of derivatives ∂̸$\not{\partial}_{a}∂$ defined for spinors and Dirac adjoint spinors through
∂̸$\not{\partial}_{a}\psi∂$ $=$ $\gamma^{a}\partial_{a}\psi~{}~{}~{}\left(a!\text{, thus no summation over }a\right)~{},$
∂̸$\not{\partial}_{a}\bar{\psi}∂$ $=$ $\partial_{a}\bar{\psi}\gamma^{a}~{}~{}~{}\left(a!\right)~{},$ (11)
respectively. These will be useful for separating temporal and spatial derivatives, as required in a Hamiltonian treatment. The ∂̸$\not{\partial}_{a}∂$ derivatives exhibit all usual properties of derivative operators, including the Leibniz rule, also the property
∂̸$\not{\partial}=\sum_{a=0}^{3}\not{\partial}_{a}~{}.∂$ (12)
Hence all suitable Lagrangians can be rewritten as $\mathfrak{L}\left(\psi,\not{\partial}_{a}\psi,\bar{\psi},\not{\partial}_{a}\bar{\psi}\right)$. By carrying out partial integrations in the action, the alternative form of the Euler–Lagrange equations emerge:
$\frac{\partial\mathfrak{L}}{\partial\psi}-\not{\partial}_{a}\frac{\partial\mathfrak{L}}{\partial\not{\partial}_{a}\psi}=0~{},\qquad\frac{\partial\mathfrak{L}}{\partial\bar{\psi}}-\not{\partial}_{a}\frac{\partial\mathfrak{L}}{\partial\not{\partial}_{a}\bar{\psi}}=0~{}.$ (13)
(The Einstein summation convention applies to the indices of the ∂̸$\not{\partial}_{a}∂$ derivatives.) In detail, they read:
$\frac{\partial\mathfrak{L}}{\partial\psi}-\sum_{a=0}^{3}\partial_{a}\frac{\partial\mathfrak{L}}{\partial\left(\gamma^{a}\partial_{a}\psi\right)}\gamma^{a}=0~{},\qquad\frac{\partial\mathfrak{L}}{\partial\bar{\psi}}-\sum_{a=0}^{3}\gamma^{a}\partial_{a}\frac{\partial\mathfrak{L}}{\partial\left(\partial_{a}\bar{\psi}\gamma^{a}\right)}=0~{}.$ (14)
(Due to the triple occurrence of indices, here the summations need to be explicitly indicated.) As $\left(\gamma^{0}\right)^{2}=\mathrm{I}_{4}$ and $\left(\gamma^{\mu}\right)^{2}=-\mathrm{I}_{4}$ the Euler–Lagrange equations (14) formally reduce to (10).
In what follows, we will employ the ∂̸$\not{\partial}_{a}∂$ derivatives instead of partial derivatives, as we found them extremely advantageous for simplifying the formalism. For example, the ∂̸$\not{\partial}_{a}∂$ derivatives of spinors (or Dirac adjoint spinors) are also spinors (or Dirac adjoint spinors). In addition, the derivatives of $\mathfrak{L}$ with respect to ∂̸$\not{\partial}_{a}\psi∂$ and ∂̸$\not{\partial}_{a}\bar{\psi}∂$ become Dirac adjoint and Dirac spinors, respectively (a property not shared by the derivatives with respect to $\partial_{a}\psi$ and $\partial_{a}\bar{\psi}$).
II.3 The simplest Lagrangian
The Lagrangian generating the Dirac equation (1) presented in many quantum field theory textbooks LancasterBlundell ; BD ; LL ; HK is
$\mathfrak{L}_{BD}=\bar{\psi}\left(i\hbar c\gamma^{a}\partial_{a}-mc^{2}\right)\psi~{},$ (15)
[…]
$\mathfrak{L}_{BD}=\bar{\psi}\left(i\hbar c\not{\partial}-mc^{2}\right)\psi~{}.$ (16)
Variation of the action with respect to $\bar{\psi}$ (or equivalently, inserting $\mathfrak{L}_{BD}$ into the Euler–Lagrange equations (13)) immediately gives the Dirac equation (1) in the form
$\left(i\frac{\hbar}{c}\not{\partial}-m\right)\psi=0~{},$ (17)
while varying with respect to $\psi$ leads to
$\left(i\frac{\hbar}{c}\not{\partial}+m\right)\bar{\psi}=0~{},$ (18)
which, by virtue of Eqs. (8) is the Dirac adjoint of the Dirac equation.
The Bjorken–Drell Lagrangian $\mathfrak{L}_{BD}$ is analogue to the one given by Schiff for the Schrödinger field, which has been discussed in Ref. Schrod from a variational point of view. We define the momenta as
$\pi^{BD}$ $=\frac{\partial\mathfrak{L}_{BD}}{\partial\not{\partial}_{0}\psi}=i\hbar c\bar{\psi}~{},$
$\bar{\pi}^{BD}$ $=\frac{\partial\mathfrak{L}_{BD}}{\partial\not{\partial}_{0}\bar{\psi}}=0~{}.$ (19)
This definition differs from the one explored in the literature by containing the ∂̸$\not{\partial}_{0}∂$ derivative, resulting in $\pi^{BD}$ without a $\gamma^{0}$ multiplier as compared to Ref. BD . This choice however conveniently ensures that $\pi^{BD}$ is a Dirac adjoint spinor. Note that momenta with an overbar are spinors and those without are Dirac adjoint spinors (opposite convention as for the fields).
This modification of their definition does not cure the problem that the canonical momenta of the Dirac adjoint variables $\psi$ and $\bar{\psi}$ fail to be the Dirac adjoints of each other (as $\bar{\pi}^{BD}$ is vanishing). A similar disadvantageous property was exhibited in the canonical treatment of the Schrödinger field by Schiff Schrod .
Eliminating $\bar{\psi}$ from the action through the first relation (19) leads to the action in already Hamiltonian form
$S_{BD}\left[\psi,\pi^{BD}\right]=\int\mathrm{d}t\int\mathrm{d}\mathbf{r}~{}\left(\pi^{BD}\not{\partial}_{0}\psi-\mathcal{H}_{BD}\right)~{},$ (20)
with the Hamiltonian
$\mathcal{H}_{BD}=-\pi^{BD}\sum_{\mu=1}^{3}\not{\partial}_{\mu}\psi-\frac{i}{\hbar}mc\left(\pi^{BD}\psi\right)~{}.$ (21)
This Hamiltonian is expressed in terms of a single canonical pair of spinors $\left(\psi,\pi^{BD}\right)$. The canonical equation emerging from the variation of the functional (20) with respect to $\pi^{BD}$ gives the Dirac equation
$\frac{\delta S_{BD}\left[\psi,\pi^{BD}\right]}{\delta\pi^{BD}}=\left(\not{\partial}+\frac{i}{\hbar}mc\right)\psi=0~{},$ (22)
while the variation with respect to $\psi$ yields
$\frac{\delta S_{BD}\left[\psi,\pi^{BD}\right]}{\delta\psi}=\left(-\not{\partial}+\frac{i}{\hbar}mc\right)\pi^{BD}=0~{},$ (23)
which, after reintroducing $\bar{\psi}$ in place of $\pi^{BD}$ is but the Dirac adjoint Dirac equation (18). Hence, there is an obvious redundancy in this variational approach.
Refs. LancasterBlundell ; BD advocate for introducing the Dirac equation into the Hamiltonian $\mathcal{H}_{BD}$ before second quantization, a procedure yielding
$\mathcal{H}_{BD2}=\pi^{BD}\not{\partial}_{0}\psi~{}.$ (24)
This however, destroys the variational principle, as it leads to a vanishing action through Eq. (20), from which no Dirac equation can be derived.
II.4 Hermitian Lagrangian
Instead of the Hamiltonian (21), expressed in terms of a single canonical pair of spinors $\left(\psi,\pi^{BD}\right)$, another Hamiltonian of two canonical pairs of complex spinors arises from the Hermitian Lagrangian IZ ; LL
$\mathfrak{L}_{IZ}=\frac{i\hbar c}{2}\left[\bar{\psi}\gamma^{a}\left(\partial_{a}\psi\right)-\left(\partial_{a}\bar{\psi}\right)\gamma^{a}\psi\right]-mc^{2}\bar{\psi}\psi$ (25)
(where the subscript IZ refers to Itzykson and Zuber). This is the analogue of the Henley–Thirring Lagrangian for the Schrödinger field, discussed from a variational point of view in Ref. Schrod , which also leads to a Hamiltonian in terms of two complex fields. The Lagrangian $\mathfrak{L}_{IZ}$ arises from $\mathfrak{L}_{BD}$ by subtracting a total divergence $\partial_{a}\left[\left(i\hbar c/2\right)\bar{\psi}\gamma^{a}\psi\right]=\left(i\hbar c/2\right)\not{\partial}\left(\bar{\psi}\psi\right)$.
[…]
$\mathfrak{L}_{IZ}=\frac{i\hbar c}{2}\left(\bar{\psi}\not{\partial}\psi-\not{\partial}\bar{\psi}\psi\right)-mc^{2}\bar{\psi}\psi$ (26)
Variation with respect to the pair of independent spinor variables $\bar{\psi}$ and $\psi$ give the Dirac equation (17) and its Dirac adjoint (18), respectively. The canonical momenta emerge as
$\pi$ $=\frac{\partial\mathfrak{L}_{IZ}}{\partial\not{\partial}_{0}\psi}=\frac{i\hbar c}{2}\bar{\psi}~{},$
$\bar{\pi}$ $=\frac{\partial\mathfrak{L}_{IZ}}{\partial\not{\partial}_{0}\bar{\psi}}=-\frac{i\hbar c}{2}\psi~{}.$ (27)
They are each other’s Dirac adjoint spinors. As before, $\pi$ is a Dirac adjoint spinor (different by a factor of $\gamma^{0}$ from the respective expression in the literature), while $\bar{\pi}$ is a spinor.
Similarly to the case of the Henley–Thirring Lagrangian for the Schrödinger field Schrod , here too we can arrive to an already Hamiltonian form of the Lagrangian $\mathfrak{L}_{IZ}$ through the following recipe. First in the kinetic term we replace all $\psi$ and $\bar{\psi}$ appearing algebraically with their expressions derived from Eqs. (27). Then we separate the mass term into two equal contributions, replacing $\psi$ in one of the halves and $\bar{\psi}$ in the other one with the respective expressions containing the momenta. In this way we arrive to
$\mathfrak{L}_{IZ}=\pi\not{\partial}_{0}\psi+\not{\partial}_{0}\bar{\psi}\bar{\pi}-\mathcal{H}_{IZ}~{}.$ (28)
Beside the Liouville form the Hamiltonian density
$\mathcal{H}_{IZ}=-\pi\sum_{\mu=1}^{3}\not{\partial}_{\mu}\psi-\sum_{\mu=1}^{3}\not{\partial}_{\mu}\bar{\psi}\bar{\pi}-\frac{imc}{\hbar}\left(\pi\psi-\bar{\psi}\bar{\pi}\right)$ (29)
emerged. In contrast with the Hamiltonian density (21) obtained in the earlier approach, this is Hermitian. Varying the action
$S_{IZ}\left[\psi,\pi,\bar{\psi},\bar{\pi}\right]=\int\mathrm{d}t\int\mathrm{d}\mathbf{r}~{}\left[\pi\not{\partial}_{0}\psi+\not{\partial}_{0}\bar{\psi}\bar{\pi}-\mathcal{H}_{IZ}\right]~{}$ (30)
with respect to $\bar{\psi}$ or $\pi$ gives the Dirac equation (17), while the variation with respect to $\psi$ or $\bar{\pi}$ gives its Dirac adjoint (18):
$-\frac{2}{c^{2}}\frac{\delta S_{IZ}\left[\psi,\pi,\bar{\psi},\bar{\pi}\right]}{\delta\psi}$ $=\left(i\frac{\hbar}{c}\not{\partial}+m\right)\bar{\psi}=0~{},$
$i\frac{\hbar}{c}\frac{\delta S_{IZ}\left[\psi,\pi,\bar{\psi},\bar{\pi}\right]}{\delta\pi}$ $=\left(i\frac{\hbar}{c}\not{\partial}-m\right)\psi=0~{},$
$\frac{2}{c^{2}}\frac{\delta S_{IZ}\left[\psi,\pi,\bar{\psi},\bar{\pi}\right]}{\delta\bar{\psi}}$ $=\left(i\frac{\hbar}{c}\not{\partial}-m\right)\psi=0~{},$
$i\frac{\hbar}{c}\frac{\delta S_{IZ}\left[\psi,\pi,\bar{\psi},\bar{\pi}\right]}{\delta\bar{\pi}}$ $=\left(i\frac{\hbar}{c}\not{\partial}+m\right)\bar{\psi}=0~{}.$ (31)
Note that all other recipes for interchanging the pair $\left(\psi,\bar{\psi}\right)$ with corresponding expressions of $\left(\pi,\bar{\pi}\right)$ fail to generate the correct Hamiltonian and field equations. This is consequence of the heavy redundancy in the choice of variables, which leads to the desired Dirac equation in four equivalent copies.
III The Dirac–Bergmann algorithm
Similarly to the Schrödinger field, either of the Lagrangians (16) or (26) characterizing the dynamics of the Dirac spinor contain linearly the generalized velocities ∂̸$\not{\partial}_{0}\psi∂$ and ∂̸$\not{\partial}_{0}\bar{\psi}∂$, a simptom of a contrained system. Hence the Dirac–Bergmann algorithm needs to be applied for an unambiguous transition from the Lagrangian to the Hamiltonian description.
For this we start from the Hermitian Lagrangian (26). The generalized momenta (27) represent two primary Hamiltonian constraints
$\bar{\phi}$ $:=\pi-\frac{i\hbar c}{2}\bar{\psi}=0~{},$
$\phi$ $:=\bar{\pi}+\frac{i\hbar c}{2}\psi=0~{}.$ (32)
The first of this is a constraint on Dirac adjoint spinors, while the second on spinors and they are the Dirac adjoint of each other.
III.1 Primary Hamiltonian
The primary Hamiltonian density arises from the naïve Legendre transformation
$\mathcal{H}_{P}=\pi\left(\not{\partial}_{0}\psi\right)+\left(\not{\partial}_{0}\bar{\psi}\right)\bar{\pi}-\mathfrak{L}_{IZ}~{},$ (33)
which can be decomposed as
$\mathcal{H}_{P}=\mathcal{H}_{C}+\bar{\phi}\left(\not{\partial}_{0}\psi\right)+\left(\not{\partial}_{0}\bar{\psi}\right)\phi~{},$ (34)
thus as the sum of the canonical Hamiltonian density
$\mathcal{H}_{C}=\frac{i\hbar c}{2}\sum_{\mu=1}^{3}\left[\left(\not{\partial}_{\mu}\bar{\psi}\right)\psi-\bar{\psi}\left(\not{\partial}_{\mu}\psi\right)\right]+mc^{2}\bar{\psi}\psi~{}$ (35)
with a combination of the primary contraints, in which the spinor ∂̸$\not{\partial}_{0}\psi∂$ and the Dirac adjoint spinor ∂̸$\not{\partial}_{0}\bar{\psi}∂$ are unknown functions of the phase space variables.
III.2 Factor ordered Poisson bracket
We introduce a factor ordered Poisson bracket (defined by the requirement that in each term a Dirac adjungate spinor multiplies a spinor) of two phase space functions $f~{}$and $g$ (thus depending only of inner products of Dirac adjungate spinors with spinors) as
$\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}$ $=$ $\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}\right.$ (36)
$\left.+\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\bar{\pi}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\psi}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\pi}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\bar{\psi}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)~{}.$
The evolution of $f$ is generated by
∂̸$\not{\partial}_{0}f\left(t,\mathbf{r}\right)\approx\left\{f\left(t,\mathbf{r}\right),H_{P}\right\}~{},∂$ (37)
[…]
Note that here the definition of the momenta (27) pays off, as it gets rid of all $\gamma^{0}$ factors, which would otherwise appear in the Poisson bracket.
III.3 Dirac equation as consistency condition
Eq. (37) allows to check the consistency of the primary constraints (32), by imposing that they stay conserved in time:
00 0 ∂̸$\approx\not{\partial}_{0}\bar{\phi}\approx\left\{\bar{\phi},H_{P}\right\}=-\frac{\delta H_{P}}{\delta\psi}-\frac{\delta H_{P}}{\delta\bar{\pi}}\frac{i\hbar c}{2}≈ ∂$
$\approx-\frac{\delta}{\delta\psi}\int\mathrm{d}\mathbf{r~{}}\mathcal{H}_{C}-\frac{i\hbar c}{2}\frac{\delta}{\delta\bar{\pi}}\int\mathrm{d}\mathbf{r~{}}\mathcal{H}_{C}-i\hbar c\not{\partial}_{0}\bar{\psi}$
$=-c^{2}\left(i\frac{\hbar}{c}\not{\partial}\bar{\psi}+m\bar{\psi}\right)$ (38)
and
00 0 ∂̸$\approx\not{\partial}_{0}\phi\approx\left\{\phi,H_{P}\right\}=\frac{i\hbar c}{2}\frac{\delta H_{P}}{\delta\pi}-\frac{\delta H_{P}}{\delta\bar{\psi}}≈ ∂$
$\approx\frac{i\hbar c}{2}\frac{\delta}{\delta\pi}\int\mathrm{d}\mathbf{r~{}}\mathcal{H}_{C}-\frac{\delta}{\delta\bar{\psi}}\int\mathrm{d}\mathbf{r~{}}\mathcal{H}_{C}+i\hbar c\not{\partial}_{0}\psi$
$=c^{2}\left(i\frac{\hbar}{c}\not{\partial}\psi-m\psi\right)~{}.$ (39)
Note, that Eqs. (38)-(39) represent shorthand notations for the spinorial components of the constraints $\bar{\phi}$ and $\phi$ (a Dirac adjoint spinor and a spinor, respectively). This is because the Poisson bracket was defined on phase space functions.
Remarkably, the Dirac equation and its Dirac adjoint both emerged as consistency conditions, holding weakly. They can also be perceived as the equations determining the unknown phase space functions ∂̸$\not{\partial}_{0}\psi∂$ and ∂̸$\not{\partial}_{0}\bar{\psi}∂$.
III.4 Time derivative
In the slashed time derivative of a phase space function  $f$ factor ordering also needs to be imposed:
∂̸$\not{\partial}_{0}f\left(t,\mathbf{r}\right)∂$ $=$ $\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime\prime}\right)}\not{\partial}_{0}\psi\left(t,\mathbf{r}^{\prime\prime}\right)+\not{\partial}_{0}\bar{\psi}\left(t,\mathbf{r}^{\prime\prime}\right)\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\psi}\left(t,\mathbf{r}^{\prime\prime}\right)}\right.$ (40)
+∂̸$\not{\partial}_{0}\pi\left(t,\mathbf{r}^{\prime\prime}\right)\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}+\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\pi}\left(t,\mathbf{r}^{\prime\prime}\right)}\not{\partial}_{0}\bar{\pi}\left(t,\mathbf{r}^{\prime\prime}\right)\right)~{}.+ ∂$
The calculation of the factor ordered Poisson bracket (36) for $g=H_{P}$ and exploring Eqs. (32), (34) and (35) yields
$\left\{f\left(t,\mathbf{r}\right),H_{P}\right\}$ $=$ ∂̸$\not{\partial}_{0}f\left(t,\mathbf{r}\right)-\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(i\hbar c\not{\partial}+mc^{2}\right)\bar{\psi}\left(t,\mathbf{r}^{\prime\prime}\right)\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}∂$ (41)
$+\int\mathrm{d}\mathbf{r}^{\prime\prime}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\pi}\left(t,\mathbf{r}^{\prime\prime}\right)}\left(i\hbar c\not{\partial}-mc^{2}\right)\psi\left(t,\mathbf{r}^{\prime\prime}\right)~{},$
leading for the phase space variables to the brackets
$\left\{\psi\left(t,\mathbf{r}\right),H_{P}\right\}$ $=$ ∂̸$\not{\partial}_{0}\psi\left(t,\mathbf{r}\right)~{},∂$
$\left\{\bar{\psi}\left(t,\mathbf{r}\right),H_{P}\right\}$ $=$ ∂̸$\not{\partial}_{0}\bar{\psi}\left(t,\mathbf{r}\right)~{},∂$
$\left\{\pi\left(t,\mathbf{r}\right),H_{P}\right\}$ $=$ ∂̸$\not{\partial}_{0}\pi\left(t,\mathbf{r}\right)-\left(i\hbar c\not{\partial}+mc^{2}\right)\bar{\psi}\left(t,\mathbf{r}\right)~{},∂$
$\left\{\bar{\pi}\left(t,\mathbf{r}\right),H_{P}\right\}$ $=$ ∂̸$\not{\partial}_{0}\bar{\pi}\left(t,\mathbf{r}\right)+\left(i\hbar c\not{\partial}-mc^{2}\right)\psi\left(t,\mathbf{r}\right)~{}.∂$ (42)
The first two of them are manifestly of the form (37), while the third and fourth take the same form when regarded as weak equalities. Indeed, the consistency conditions of the constraints imply the Dirac equation and its Dirac adjoint.
III.5 Factor ordered Dirac bracket
We can also establish, that the constraints are second class
$\left\{\bar{\phi},\phi\right\}:=\left\{\bar{\phi}\left(t,\mathbf{r}\right),\phi\left(t,\mathbf{r}^{\prime}\right)\right\}=-i\hbar c\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (43)
This holds component-wise.
In order to rewrite time evolution in terms of a Dirac bracket, we need the inverse $A^{-1}$ of the distributional matrix
$A\left(\mathbf{r}-\mathbf{r}^{\prime}\right)=\left(\left\{\phi^{i}\left(t,\mathbf{r}\right),\phi^{j}\left(t,\mathbf{r}^{\prime}\right)\right\}\right)=\left(\begin{array}[]{cc}0&-i\hbar c\\ i\hbar c&0\end{array}\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (44)
where $\phi^{i}$ stands for $\bar{\phi}$ ($i=1$) and $\phi$ ($i=2$), defined as
[…]
which is
$A^{-1}\left(\mathbf{r}-\mathbf{r}^{\prime}\right)=\left(\begin{array}[]{cc}0&-\frac{i}{\hbar c}\\ \frac{i}{\hbar c}&0\end{array}\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (46)
We also need
$\left\{f\left(t,\mathbf{r}\right),\bar{\phi}\left(t,\mathbf{r}^{\prime}\right)\right\}$ $=$ $\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime}\right)}+\frac{i\hbar c}{2}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)}~{},$
$\left\{f\left(t,\mathbf{r}\right),\phi\left(t,\mathbf{r}^{\prime}\right)\right\}$ $=$ $-\frac{i\hbar c}{2}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime}\right)}+\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)}~{}.$ (47)
These are a Dirac adjoint spinor and a spinor, respectively. The Dirac bracket of two functions $f$, $g$ is defined as
$\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}\!$
$-\!\!\int\!\!\!\int\!\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\left\{f\left(t,\mathbf{r}\right),\phi^{i}\left(t,\mathbf{r}^{\prime\prime}\right)\right\}A_{ij}^{-1}\left(\mathbf{r}^{\prime\prime}\!-\mathbf{r}^{\prime\prime\prime}\right)\left\{\phi^{j}\left(t,\mathbf{r}^{\prime\prime\prime}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}~{}.$ (48)
In the matrix multiplication one should pay attention to the factor ordering, aiming for Dirac adjoint spinor times spinor type inner products in each term. By this we mean that the two nonvanishing terms in the sum have a different factor ordering, resulting in
$\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}$
$-\!\!\int\!\!\!\int\!\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\left\{f\left(t,\mathbf{r}\right),\bar{\phi}\left(t,\mathbf{r}^{\prime\prime}\right)\right\}A_{12}^{-1}\left(\mathbf{r}^{\prime\prime}\!-\mathbf{r}^{\prime\prime\prime}\right)\left\{\phi\left(t,\mathbf{r}^{\prime\prime\prime}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}$
$-\!\!\int\!\!\!\int\!\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\left\{\bar{\phi}\left(t,\mathbf{r}^{\prime\prime\prime}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}A_{21}^{-1}\left(\mathbf{r}^{\prime\prime}\!-\mathbf{r}^{\prime\prime\prime}\right)\left\{f\left(t,\mathbf{r}\right),\phi\left(t,\mathbf{r}^{\prime\prime}\right)\right\}~{}.$ (49)
Calculation gives:
$\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{1}{2}\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}$
$-\frac{i}{\hbar c}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\bar{\psi}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\psi}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)$
$-\frac{i\hbar c}{4}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\bar{\pi}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\bar{\pi}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}\right)~{}.$ (50)
Formally, this resembles the Dirac bracket derived for the Schrödinger equation (Eq. (3.7) of Ref. Schrod ), with the sole (but essential) difference of factor ordering.
Straightforward calculation shows, that
∂̸$\not{\partial}_{0}f\approx\left\{f,H_{P}\right\}_{D}~{}∂$ (51)
also holds, as it should for a constrained system without first class constraints.
III.6 Dirac brackets of canonical data
The Dirac brackets (50) of the canonical phase space variables emerge as
$\left\{\psi\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=-\frac{i}{\hbar c}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (52)
$\left\{\pi\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{i\hbar c}{4}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (53)
$\left\{\psi\left(t,\mathbf{r}\right),\psi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\psi}\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{}.$ (54)
$\left\{\pi\left(t,\mathbf{r}\right),\pi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\pi}\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{},$ (55)
$\left\{\psi\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\psi}\left(t,\mathbf{r}\right),\pi\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{},$ (56)
$\left\{\psi\left(t,\mathbf{r}\right),\pi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{1}{2}\left\{\psi\left(t,\mathbf{r}\right),\pi\left(t,\mathbf{r}^{\prime}\right)\right\}=\frac{1}{2}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (57)
$\left\{\bar{\psi}\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{1}{2}\left\{\bar{\psi}\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}=\frac{1}{2}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (58)
When inserting the definitions (27) of the momenta, all nontrivial Dirac brackets reduce to Eqs. (57)-(58).
III.7 Reduced phase space
The Poisson bracket (43) of the constrains generating a delta function suggests to introduce the rescaled constraints as a canonical pair as
$\left(\begin{array}[]{c}\psi_{2}=-\frac{i}{\hbar c}\phi\\ \pi_{2}=\bar{\phi}\end{array}\right)~{}.$ (59)
These indeed obey $\left\{\psi_{2}\left(t,\mathbf{r}\right),\pi_{2}\left(t,\mathbf{r}^{\prime}\right)\right\}=\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)$. The other pair of canonical variables $\left(\psi_{1},\pi_{1}\right)$ then emerges by imposing $\left\{\psi_{1}\left(t,\mathbf{r}\right),\pi_{1}\left(t,\mathbf{r}^{\prime}\right)\right\}=\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)$, $\left\{\psi_{1}\left(t,\mathbf{r}\right),\psi_{2}\left(t,\mathbf{r}^{\prime}\right)\right\}=0$, $\left\{\psi_{1}\left(t,\mathbf{r}\right),\pi_{2}\left(t,\mathbf{r}^{\prime}\right)\right\}=0$, $\left\{\psi_{2}\left(t,\mathbf{r}\right),\pi_{1}\left(t,\mathbf{r}^{\prime}\right)\right\}=0$ and $\left\{\pi_{1}\left(t,\mathbf{r}\right),\pi_{2}\left(t,\mathbf{r}^{\prime}\right)\right\}=0$ as
$\left(\begin{array}[]{c}\psi_{1}=\frac{1}{2}\psi+\frac{i}{\hbar c}\bar{\pi}\\ \pi_{1}=\frac{i\hbar c}{2}\bar{\psi}+\pi\end{array}\right)~{}.$ (60)
In the new canonical chart $\psi_{1,2}$ are spinors, while $\pi_{1,2}$ are Dirac adjoint spinors. The reduced phase space is defined as the surface of the phase space, where the constraints are obeyed, thus given as $\psi_{2}=0=\pi_{2}$ and coordinatized by $\left(\psi_{1},\pi_{1}\right)$ only.
Similarly as it happened for the Schrödinger case, here too a straightforward check shows that the Dirac bracket (50) written in terms of the new coordinates becomes the Poisson bracket of the reduced phase space. Indeed, under the canonical transformation the Poisson bracket in the variables $\left(\psi,\pi,\bar{\psi},\bar{\pi}\right)$ transforms to the Poisson bracket in the variables $\left(\psi_{1},\pi_{1},\psi_{2},\pi_{2}\right)$:
$\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}=\int\mathrm{d}\mathbf{r}^{\prime\prime}\sum_{i=1}^{2}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{i}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{i}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{i}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{i}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)~{}.$ (61)
Remarkably, in the new canonical chart there is no need for its factor ordering, as the Dirac adjoint spinor times spinor structure emerges naturally.
The complementary terms in the Dirac bracket arising from the second class constraints give
$\frac{1}{2}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)$
$-\frac{1}{2}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)$ (62)
Thus the Dirac bracket becomes
$\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)~{},$ (63)
which is exactly the Poisson bracket on the reduced phase space. It is straightforward to verify the generic property that the Dirac brackets of the constraints $\psi_{2}$ and $\pi_{2}$ with arbitrary functions vanish.
The time evolutions are still given as
∂̸$\not{\partial}_{0}f\approx\left\{f,H_{P}\right\}_{D}~{}.∂$ (64)
To prove this, we note that the primary Hamiltonian density in the new canonical chart becomes
$\mathcal{H}_{P}$ $=$ $\frac{1}{2}\left[\not{\partial}\left(\pi_{1}-\pi_{2}\right)\right]\psi_{2}+\frac{1}{2}\pi_{2}\not{\partial}\left(\psi_{1}+\psi_{2}\right)$ (65)
$+\frac{1}{2}\sum_{\mu=1}^{3}\left[\not{\partial}_{\mu}\left(\pi_{1}-\pi_{2}\right)\right]\psi_{1}+\frac{1}{2}\left[\not{\partial}_{0}\left(\pi_{1}-\pi_{2}\right)\right]\psi_{2}$
$-\frac{1}{2}\pi_{1}\sum_{\mu=1}^{3}\not{\partial}_{\mu}\left(\psi_{1}+\psi_{2}\right)+\frac{1}{2}\pi_{2}\not{\partial}_{0}\left(\psi_{1}+\psi_{2}\right)$
$-\frac{ic}{\hbar}m\left(\pi_{1}-\pi_{2}\right)\left(\psi_{1}+\psi_{2}\right)~{}.$
Then,
$\left\{f,H_{P}\right\}_{D}~{}$ $=$ $\int\mathrm{d}\mathbf{r}^{\prime\prime}\left[\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\not{\partial}_{0}\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)+\not{\partial}_{0}\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\right.$ (66)
$+\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{ic}{\hbar}\left(i\frac{\hbar}{c}\not{\partial}-m\right)\left(\psi_{1}+\psi_{2}\right)\left(t,\mathbf{r}^{\prime\prime}\right)$
$\left.+\frac{ic}{\hbar}\left(i\frac{\hbar}{c}\not{\partial}+m\right)\left(\pi_{1}-\pi_{2}\right)\left(t,\mathbf{r}^{\prime\prime}\right)\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\right]~{}.$
On the constraint surface the first line is but the slashed time derivative, as the consistency conditions can be rewritten as ∂̸$\not{\partial}_{0}\psi_{2}=0=\not{\partial}_{0}\pi_{2}∂$, allowing to add the two missing terms to the total time derivative. The second and third lines also vanish on shell as $\psi_{1}+\psi_{2}=\psi$ and $\pi_{1}-\pi_{2}=i\hbar c\bar{\psi}$ (which obey the Dirac equation and its Dirac adjoint, as consistency conditions).
We conclude the subsection by noting that the Hamiltonian $\mathcal{H}_{R}$ on the reduced phase space, obtained by rewriting $\mathcal{H}_{P}\approx\mathcal{H}_{C}$ in the canonical chart $\left(\psi_{1},\pi_{1},\psi_{2}=0,\pi_{2}=0\right)$ can be expressed in the alternative forms
$\mathcal{H}_{R}$ $=$ $\frac{1}{2}\sum_{\mu=1}^{3}\left[\left(\not{\partial}_{\mu}\pi_{1}\right)\psi_{1}-\pi_{1}\left(\not{\partial}_{\mu}\psi_{1}\right)\right]-\frac{i}{\hbar}mc\pi_{1}\psi_{1}$ (67)
$=$ $\sum_{\mu=1}^{3}\not{\partial}_{\mu}\left(\frac{\pi_{1}\psi_{1}}{2}\right)+\mathcal{H}_{BD}\left(\psi_{1},\pi_{1}\right)$
$=$ $\mathcal{H}_{IZ}\left(\psi_{1},\pi_{1},\psi_{2}=0,\pi_{2}=0\right)~{}.$
This is slightly different from the Schrödinger case, where the Hamiltonian on the reduced phase space coincides with both the Schiff and Henley–Thirring Hamiltonians, when evaluated on shell.
Variation of the action
$S_{R}=\int\mathrm{d}^{4}x\left[\pi_{1}\not{\partial}_{0}\psi_{1}-\mathcal{H}_{R}\right]$ (68)
with respect to the reduced space variables $\psi_{1}$ and $\pi_{1}$ gives
$\frac{\delta S_{R}\left[\psi_{1},\pi_{1}\right]}{\delta\psi_{1}}$ $=i\frac{c}{\hbar}\left(i\frac{\hbar}{c}\not{\partial}+m\right)\pi_{1}=0~{},$
$\frac{\delta S_{R}\left[\psi_{1},\pi_{1}\right]}{\delta\pi_{1}}$ $=-i\frac{c}{\hbar}\left(i\frac{\hbar}{c}\not{\partial}-m\right)\psi_{1}=0~{}.$ (69)
These are the Dirac equation and its Dirac adjoint, respectively, both obtained as canonical equations on the reduced phase space.
Hence, at a classical level the Dirac field can be suitably characterized through a Hamiltonian evolution through the Dirac–Bergmann method, in full analogy with the corresponding procedure for the Schrödinger field. The price to pay was the need for introducing factor ordered Poisson and Dirac brackets. The discussion was simplified by the modified definitions of the momenta as compared to the standard literature. Then we gave the Dirac brackets of the canonical data as Eqs. (54)-(58). We also identified the reduced phase space and the reduced Hamiltonian (67) on it.
IV The Dirac field as Grassmann variable
In an alternative approach explored in Refs. HT and Grassmann2 , classically the Dirac field is regarded as an odd Grassmann variable. For comparison we explicitly present the Dirac–Bergmann algorithm for the Dirac field, treated as an odd Grassmann variable, in two equivalent formulations. For simplifying the notation, we also drop the arguments of the variables in this section.
IV.1 Grassmann variables
The generators $\xi^{i}$ of a Grassmann algebra anticommute as
$\xi^{i}\xi^{j}=-\xi^{j}\xi^{i}~{}.$ (70)
In the above notation we have supressed the exterior or wedge product between them. Such products applied repeatedly on the generators lead to superfunctions. By definition, a subclass of superfunctions anticommuting with each other are odd (of parity $1$). Another subclass, consisting of superfunctions commuting with each other and also with the odd ones are even (of parity 00). The simplest superfunctions are the odd (even) Grassmann variables, also dubbed Grassmann numbers, defined as series of odd (even) number products of generators with complex coefficients. Classically the Dirac spinor, its adjoint and Dirac adjoint all qualify as odd Grassman variables, which anticommute, for example $\bar{\psi}\psi=-\psi\bar{\psi}$. By contrast, an even Grassmann variable $E$ commutes with both even or odd Grassman variables, or in fact with any superfunction $F$ as $EF=FE$. Following this approach there is no need for factor ordering, as $\psi,\ \psi^{{\dagger}},\ \bar{\psi}$ appear on equal footing (they can be thought of as $4\times 4$ matrices). The price to pay is that when calculating the brackets of the superfunctions, two versions of partial derivatives with respect to the generators need to be introduced. These are the left (denoted $L$, acting from the left to the right) and right ($R$, acting from the right to the left) derivatives, related as
$\dfrac{\partial^{L}F}{\partial\xi^{i}}=(-1)^{1+\epsilon_{F}}\dfrac{\partial^{R}F}{\partial\xi^{i}}~{},$ (71)
where $\epsilon_{F}$ indicates the parity of $F$. We illustrate the action of these derivatives on the product $\xi^{1}\xi^{2}$ as follows
$\dfrac{\partial^{L}}{\partial\xi^{1}}\left(\xi^{1}\xi^{2}\right)$ $=$ $\xi^{2}~{},\quad\dfrac{\partial^{L}}{\partial\xi^{2}}\left(\xi^{1}\xi^{2}\right)=-\xi^{1}~{},$
$\dfrac{\partial^{R}}{\partial\xi^{1}}\left(\xi^{1}\xi^{2}\right)$ $=$ $-\xi^{2}~{},\quad\dfrac{\partial^{R}}{\partial\xi^{2}}\left(\xi^{1}\xi^{2}\right)=\xi^{1}~{}.$ (72)
In general, the action of a left derivative on a product of any Grassmann superfunctions $FG$ with respect to an odd Grassmann variable $\psi$ is Berezin :
$\dfrac{\partial^{L}\left(FG\right)}{\text{ }\partial\psi}=\dfrac{\partial^{L}F}{\text{ }\partial\psi}G+(-1)^{\epsilon_{F}}F\dfrac{\partial^{L}G}{\text{ }\partial\psi}~{},$ (73)
whereas a right derivative acts as
$\dfrac{\partial^{R}\left(FG\right)}{\text{ }\partial\psi}=F\dfrac{\partial^{R}G}{\text{ }\partial\psi}+(-1)^{\epsilon_{G}}\dfrac{\partial^{R}F}{\text{ }\partial\psi}G~{}.$ (74)
Similar rules apply for differentiation with respect to $\bar{\psi}$. For the derivative with respect to an even Grassmann variable the Leibniz rule applies for both types of derivatives.
IV.2 Hamiltonian formalism in terms of left derivatives
For the Lagrangian of the Dirac system we once again adopt (26), denoted $\mathfrak{L}_{G}$ to emphasize that this time it contains Grassmann variables, hence it is a superdensity, of parity 00. As before, we define the momenta as derivatives with respect to the ∂̸$\not{\partial}_{0}∂$ derivatives of the field and its Dirac adjoint. However it has to be specified whether the derivative is left or right. While both conventions work, in this subsection we follow Ref. HT , in adopting the left derivative. We define the general momenta in terms of left derivatives as:
$\pi_{L}$ $=$ $\dfrac{\partial^{L}\mathfrak{L}_{G}}{\partial\not{\partial}_{0}\psi}=-\dfrac{i\hbar c}{2}\bar{\psi}~{},$
$\bar{\pi}_{L}$ $=$ $\dfrac{\partial^{L}\mathfrak{L}_{G}}{\partial\not{\partial}_{0}\bar{\psi}}=-\dfrac{i\hbar c}{2}\psi~{}.$ (75)
They differ by $\gamma^{0}$ factors in the denominator from the definitions advanced in Ref. HT (which contain $\partial_{0}$ instead ∂̸$\not{\partial}_{0}∂$). The purpose of this change is to remain as close as possible to the expressions of the momenta introduced in the previous chapter, Eqs. (27). Indeed, $\bar{\pi}_{L}$ has the same form as $\bar{\pi}$, however $\pi_{L}$ differs in a minus sign. The momenta are odd variables.
Following the same procedure as before (also paying attention that when employing left derivatives, in order to reproduce the momenta, the Liouville form needs to be defined as ∂̸$\not{\partial}_{0}\psi\pi_{L}+\not{\partial}_{0}\bar{\psi}\bar{\pi}_{L}∂$), the Hamilton superdensity emerges:
$\mathcal{H}_{L}=-\sum_{\mu}\not{\partial}_{\mu}\psi\pi_{L}-\sum_{\mu}\not{\partial}_{\mu}\bar{\psi}\bar{\pi}_{L}+\dfrac{imc}{\hbar}(\psi\pi_{L}+\bar{\psi}\bar{\pi}_{L})~{}.$ (76)
The Hamiltonian equations are
∂̸$\not{\partial}_{0}\psi∂$ $=$ $-\dfrac{\partial^{L}\mathcal{H}_{L}}{\partial\pi_{L}}~{},\quad\not{\partial}_{0}\bar{\psi}=-\dfrac{\partial^{L}\mathcal{H}_{L}}{\partial\bar{\pi}_{L}}~{},$
∂̸$\not{\partial}_{0}\pi_{L}∂$ $=$ $-\dfrac{\partial^{L}\mathcal{H}_{L}}{\partial\psi}~{},\quad\not{\partial}_{0}\bar{\pi}_{L}=-\dfrac{\partial^{L}\mathcal{H}_{L}}{\partial\bar{\psi}}~{}.$ (77)
After accounting for the change in the definitions of momenta, these agree with the corresponding equations presented in Ref. HT . Note that the left derivatives above can also be rewritten as right derivatives through Eq. (71).
The generalized Poisson bracket of generic superfunctions $F,G$ with respect to canonical data of the Dirac field is:
$\{F,G\}_{+}^{L}=\int\mathrm{d}\mathbf{r}\Bigg{[}(-1)^{\epsilon_{F}}\bigg{(}\dfrac{\delta^{L}F}{\delta\psi}\dfrac{\delta^{L}G}{\delta\pi_{L}}+\dfrac{\delta^{L}F}{\delta\pi_{L}}\dfrac{\delta^{L}G}{\delta\psi}+\dfrac{\delta^{L}F}{\delta\bar{\psi}}\dfrac{\delta^{L}G}{\delta\bar{\pi}_{L}}+\dfrac{\delta^{L}F}{\delta\bar{\pi}_{L}}\dfrac{\delta^{L}G}{\delta\bar{\psi}}\bigg{)}\Bigg{]}~{}.$ (78)
It obeys
$\{F,G\}_{+}^{L}=-(-1)^{\epsilon_{F}\epsilon_{G}}\{G,F\}_{+}^{L}~{}.$ (79)
The Poisson brackets of canonical variables become:
$\{\psi,\bar{\psi}\}_{+}^{L}$ $=$ $\{\bar{\psi},\psi\}_{+}^{L}=\{\pi_{L},\bar{\pi}_{L}\}_{+}^{L}=\{\bar{\pi}_{L},\pi_{L}\}_{+}^{L}=0~{},$
$\{\psi,\pi_{L}\}_{+}^{L}$ $=$ $\{\pi_{L},\psi\}_{+}^{L}=\{\bar{\psi},\bar{\pi}_{L}\}_{+}^{L}=\{\bar{\pi}_{L},\bar{\psi}\}_{+}^{L}=-\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})~{}.$ (80)
These Poisson brackets are symmetric.
IV.3 Hamiltonian formalism in terms of right derivatives
The dynamics can equally be expressed in terms of right derivatives, as follows. The momenta in this case are conveniently defined as
$\pi_{R}$ $=$ $\dfrac{\partial^{R}\mathfrak{L}_{G}}{\partial\not{\partial}_{0}\psi}=\dfrac{i\hbar c}{2}\bar{\psi}~{},$
$\bar{\pi}_{R}$ $=$ $\dfrac{\partial^{R}\mathfrak{L}_{G}}{\partial\not{\partial}_{0}\bar{\psi}}=\dfrac{i\hbar c}{2}\psi~{}.$ (81)
Note the differences in signs as compared with $\pi_{L}$ and $\bar{\pi}_{L}$.
The Liouville form now becomes $\pi_{R}\not{\partial}_{0}\psi+\bar{\pi}_{R}\not{\partial}_{0}\bar{\psi}$ in order to be in agreement with the definitions (81), leading to the Hamiltonian superdensity:
$\mathcal{H}_{R}=-\pi_{R}\sum_{\mu}\not{\partial}_{\mu}\psi-\bar{\pi}_{R}\sum_{\mu}\not{\partial}_{\mu}\bar{\psi}+\dfrac{imc}{\hbar}(\pi_{R}\psi+\bar{\pi}_{R}\bar{\psi})~{},$ (82)
an expression different from (76) in its factor ordering. The Hamiltonian equations read
∂̸$\not{\partial}_{0}\psi∂$ $=$ $-\dfrac{\partial^{R}\mathcal{H}_{R}}{\partial\pi_{R}}~{},\quad\not{\partial}_{0}\bar{\psi}=-\dfrac{\partial^{R}\mathcal{H}_{R}}{\partial\bar{\pi}_{R}}~{},$
∂̸$\not{\partial}_{0}\pi_{R}∂$ $=$ $-\dfrac{\partial^{R}\mathcal{H}_{R}}{\partial\psi}~{},\quad\not{\partial}_{0}\bar{\pi}_{R}=-\dfrac{\partial^{R}\mathcal{H}_{R}}{\partial\bar{\psi}}~{}.$ (83)
After accounting for the change in the definitions of momenta, these agree with the corresponding equations presented in Ref. Grassmann2 . They can also be rewritten in terms of left derivatives, if needed.
The generalized Poisson bracket with right derivatives is:
$\{F,G\}_{+}^{R}=\int\mathrm{d}\mathbf{r}\Bigg{[}(-1)^{\epsilon_{G}}\bigg{(}\dfrac{\delta^{R}F}{\delta\psi}\dfrac{\delta^{R}G}{\delta\pi_{R}}+\dfrac{\delta^{R}F}{\delta\pi_{R}}\dfrac{\delta^{R}G}{\delta\psi}+\dfrac{\delta^{R}F}{\delta\bar{\psi}}\dfrac{\delta^{R}G}{\delta\bar{\pi}_{R}}+\dfrac{\delta^{R}F}{\delta\bar{\pi}_{R}}\dfrac{\delta^{R}G}{\delta\bar{\psi}}\bigg{)}\Bigg{]}.$ (84)
It also obeys a similar property to (79). Finally, the Poisson brackets of canonical variables are
$\{\psi,\bar{\psi}\}_{+}^{R}$ $=$ $\{\bar{\psi},\psi\}_{+}^{R}=\{\pi_{R},\bar{\pi}_{R}\}_{+}^{R}=\{\bar{\pi}_{R},\pi_{R}\}_{+}^{R}=0~{},$
$\{\psi,\pi_{R}\}_{+}^{R}$ $=$ $\{\pi_{R},\psi\}_{+}^{R}=\{\bar{\psi},\bar{\pi}_{R}\}_{+}^{R}=\{\bar{\pi}_{R},\bar{\psi}\}_{+}^{R}=-\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})~{}.$ (85)
We remark that despite the different brackets employed here and in the previous subsection, the two algebras are identical.
IV.4 Dirac-Bergmann algorithm
For brevity we present the Dirac–Bergmann algorithm in terms of the two types of derivatives in parallel. The constrains are
$\phi^{L}=\bar{\pi}_{L}+\dfrac{i\hbar c}{2}\psi~{},\quad\bar{\phi}^{L}=\pi_{L}+\dfrac{i\hbar c}{2}\bar{\psi}~{},$ (86)
$\phi^{R}=\bar{\pi}_{R}-\dfrac{i\hbar c}{2}\psi~{},\quad\bar{\phi}^{R}=\pi_{R}-\dfrac{i\hbar c}{2}\bar{\psi}~{}.$ (87)
The primary Hamiltonian superdensities follow:
$\mathcal{H}_{PL}$ $=$ $\mathcal{H}_{L}+\not{\partial}_{0}\psi\bar{\phi}^{L}+\not{\partial}_{0}\bar{\psi}\phi^{L}$ (88)
$=$ $-\sum_{\mu}\not{\partial}_{\mu}\psi\pi_{L}-\sum_{\mu}\not{\partial}_{\mu}\bar{\psi}\bar{\pi}_{L}+\dfrac{imc}{\hbar}(\psi\pi_{L}+\bar{\psi}\bar{\pi}_{L})+\not{\partial}_{0}\psi\bar{\phi}^{L}+\not{\partial}_{0}\bar{\psi}\phi^{L}~{},$
$\mathcal{H}_{PR}$ $=$ $\mathcal{H}_{R}+\bar{\phi}^{R}\not{\partial}_{0}\psi+\phi^{R}\not{\partial}_{0}\bar{\psi}$ (89)
$=$ $-\pi_{R}\sum_{\mu}\not{\partial}_{\mu}\psi-\bar{\pi}_{R}\sum_{\mu}\not{\partial}_{\mu}\bar{\psi}+\dfrac{imc}{\hbar}(\pi_{R}\psi+\bar{\pi}_{R}\bar{\psi})+\bar{\phi}^{R}\not{\partial}_{0}\psi+\phi^{R}\not{\partial}_{0}\bar{\psi}~{}.$
As before, the time evolutions
[…]
conserve the constraints (where $\phi$ stands for any of $\phi^{L,R},~{}\bar{\phi}^{L,R}$).
The Poisson brackets of constrains:
$\{\bar{\phi}^{L},\phi^{L}\}_{+}^{L}=-\int\mathrm{d}\mathbf{r}\bigg{[}\dfrac{\delta^{L}\bar{\phi}^{L}}{\delta\pi_{L}}\dfrac{\delta^{L}\phi^{L}}{\delta\psi}+\dfrac{\delta^{L}\bar{\phi}^{L}}{\delta\bar{\psi}}\dfrac{\delta^{L}\phi^{L}}{\delta\bar{\pi}_{L}}\bigg{]}=-i\hbar c\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ ,$ (91)
$\{\bar{\phi}^{R},\phi^{R}\}_{+}^{R}=-\int\mathrm{d}\mathbf{r}\bigg{[}\dfrac{\delta^{R}\bar{\phi}^{R}}{\delta\pi_{R}}\dfrac{\delta^{R}\phi^{R}}{\delta\psi}+\dfrac{\delta^{R}\bar{\phi}^{R}}{\delta\bar{\psi}}\dfrac{\delta^{R}\phi^{R}}{\delta\bar{\pi}_{R}}\bigg{]}=i\hbar c\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})$ (92)
generate the matrices
$A_{ij}^{L}=-A_{ij}^{R}=\begin{bmatrix}0&-i\hbar c\\ -i\hbar c&0\end{bmatrix}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})~{},\qquad(A_{ij}^{L})^{-1}=-(A_{ij}^{R})^{-1}=\begin{bmatrix}0&\frac{i}{\hbar c}\\ \frac{i}{\hbar c}&0\end{bmatrix}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})~{},$ (93)
allowing to define the Dirac brackets through
$\{F,G\}_{D+}^{L,R}=\{F,G\}_{+}^{L,R}-\int\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\{f,\phi^{i}\}_{+}^{L,R}(A_{ij}^{L,R})^{-1}\{\phi^{j},g\}_{+}^{L,R}~{}.$ (94)
As a result, the following two Dirac brackets emerge
$\{F,G\}_{D+}^{L}=\dfrac{1}{2}\{F,G\}_{+}^{L}$ $+(-1)^{\epsilon_{f}}\dfrac{i}{\hbar c}\int\mathrm{d}\mathbf{r}\bigg{[}\dfrac{\delta^{L}F}{\delta\psi}\dfrac{\delta^{L}G}{\delta\bar{\psi}}+\dfrac{\delta^{L}F}{\delta\bar{\psi}}\dfrac{\delta^{L}G}{\delta\psi}\bigg{]}$
$-(-1)^{\epsilon_{F}}\dfrac{i\hbar c}{4}\int\mathrm{d}\mathbf{r}\bigg{[}\dfrac{\delta^{L}F}{\delta\pi_{L}}\dfrac{\delta^{L}G}{\delta\bar{\pi}_{L}}+\dfrac{\delta^{L}F}{\delta\bar{\pi}_{L}}\dfrac{\delta^{L}G}{\delta\pi_{L}}\bigg{]}\ ,$ (95)
$\{F,G\}_{D+}^{R}=\dfrac{1}{2}\{F,G\}_{+}^{R}$ $-(-1)^{\epsilon_{G}}\dfrac{i}{\hbar c}\int\mathrm{d}\mathbf{r}\bigg{[}\dfrac{\delta^{R}F}{\delta\psi}\dfrac{\delta^{R}G}{\delta\bar{\psi}}+\dfrac{\delta^{R}F}{\delta\bar{\psi}}\dfrac{\delta^{R}G}{\delta\psi}\bigg{]}$
$+(-1)^{\epsilon_{G}}\dfrac{i\hbar c}{4}\int\mathrm{d}\mathbf{r}\bigg{[}\dfrac{\delta^{R}F}{\delta\pi_{R}}\dfrac{\delta^{R}G}{\delta\bar{\pi}_{R}}+\dfrac{\delta^{R}F}{\delta\bar{\pi}_{R}}\dfrac{\delta^{R}G}{\delta\pi_{R}}\bigg{]}\ .$ (96)
In particular, the nontrivial Dirac brackets of the canonical variables become:
$\{\psi,\bar{\psi}\}_{D+}^{L}$ $=$ $\{\bar{\psi},\psi\}_{D+}^{L}=-\dfrac{i}{\hbar c}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ ,$
$\{\pi_{L},\bar{\pi}_{L}\}_{D+}^{L}$ $=$ $\{\bar{\pi}_{L},\pi_{L}\}_{D+}^{L}=\dfrac{i\hbar c}{4}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ ,$
$\{\psi,\pi_{L}\}_{D+}^{L}$ $=$ $\{\pi_{L},\psi\}_{D+}^{L}=\{\bar{\psi},\bar{\pi}_{L}\}_{D+}^{L}=\{\bar{\pi}_{L},\bar{\psi}\}_{D+}^{L}=-\dfrac{1}{2}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ ,$ (97)
and
$\{\psi,\bar{\psi}\}_{D+}^{R}$ $=$ $\{\bar{\psi},\psi\}_{D+}^{R}=\dfrac{i}{\hbar c}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ ,$
$\{\pi_{R},\bar{\pi}_{R}\}_{D+}^{R}$ $=$ $\{\bar{\pi}_{R},\pi_{R}\}_{D+}^{R}=-\dfrac{i\hbar c}{4}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ ,$
$\{\psi,\pi_{R}\}_{D+}^{R}$ $=$ $\{\pi_{R},\psi\}_{D+}^{R}=\{\bar{\psi},\bar{\pi}_{R}\}_{D+}^{R}=\{\bar{\pi}_{R},\bar{\psi}\}_{D+}^{R}=-\dfrac{1}{2}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ .$ (98)
Note that now only the Dirac brackets of the canonical pairs are identical when calculated in terms of left or right derivatives. Hence, the $L,R$ superscripts may be omitted from both the Poisson and Dirac brackets of the canonical pairs only. Also note the appearance of the $1/2$ factors, similarly to the Dirac brackets of the canonical pairs in the spinor formalism.
IV.5 Reduced phase space of Grassmann variables
It is easy to identify a canonical chart span by the canonical pair
$\left(\begin{array}[]{c}\psi_{1}^{L}=\frac{1}{2}\psi+\frac{i}{\hbar c}\bar{\pi}_{L}\\ \pi_{1}^{L}=-\frac{i\hbar c}{2}\bar{\psi}+\pi_{L}\end{array}\right)~{}$ (99)
on the reduced phase space and a second canonical pair
$\left(\begin{array}[]{c}\psi_{2}^{L}=\frac{1}{2}\psi+\frac{i}{\hbar c}\bar{\pi}_{L}=-\dfrac{i}{\hbar c}\phi^{L}\\ \pi_{2}^{L}=\frac{i\hbar c}{2}\bar{\psi}+\pi_{L}=\bar{\phi}^{L}\end{array}\right)~{}$ (100)
representing the constraints. We have checked that in terms of this canonical chart the Dirac bracket (95) on the full phase space becomes the Poisson bracket (78) on the reduced phase space (thus with variables $\psi_{1}^{L},\pi_{1}^{L}$ only).
Similarly the canonical pair
$\left(\begin{array}[]{c}\psi_{1}^{R}=\frac{1}{2}\psi-\frac{i}{\hbar c}\bar{\pi}_{R}\\ \pi_{1}^{R}=\frac{i\hbar c}{2}\bar{\psi}+\pi_{R}\end{array}\right)$ (101)
on the reduced phase space and the pair of contraints
$\left(\begin{array}[]{c}\psi_{2}^{R}=\frac{1}{2}\psi+\frac{i}{\hbar c}\bar{\pi}_{R}=\dfrac{i}{\hbar c}\phi^{R}\\ \pi_{2}^{R}=\pi_{R}-\frac{i\hbar c}{2}\bar{\psi}=\bar{\phi}^{R}\end{array}\right)$ (102)
generate another canonical chart for which the Dirac bracket (96) on the full phase space becomes the Poisson bracket (84) on the reduced phase space.
V Canonical Second Quantization of the Dirac field
We have seen that a classical Hamiltonian constrained treatment of the Dirac field is possible is several ways. We have shown, how a Poisson structure can be employed at the price of redefining both the Poisson and Dirac brackets such that a factor ordering necessary for the Dirac spinor interpretation of the Dirac field can be retained. Alternatively, regarding the Dirac field as a Grassman odd variable, the Hamiltonian treatment can also be pursuit in terms of generalized Poisson and generalized Dirac structures for anticommuting fields.
Heuristically, the canonical quantization can be achieved from both type of approaches, as presented below.
V.1 Spinorial approach
Due to the factor ordering requirements imposed in the classical discussion one has to pay special attention to reflect this in the canonical quantization as well.
A suitable recipe to be applied here is to transform the Dirac bracket of a Dirac spinor $\mathcal{D}$ (depending on the phase space variables $\psi,\bar{\pi}$) with a Dirac adjoint spinor $\mathcal{A}$ (depending on the phase space variables $\bar{\psi},\pi$) into an anticommutator as follows
$\left\{\mathcal{D},\mathcal{A}\right\}_{D}\mapsto-\dfrac{i}{\hbar c}[\widehat{\mathcal{D}},\widehat{\mathcal{A}}\gamma^{0}]_{+}~{}.$ (103)
In particular, for the nonvanishing Dirac brackets (52), (53), (57), (58) the above prescription yields
$\left[\widehat{\psi}\left(t,\mathbf{r}\right),\widehat{\psi}^{{\dagger}}\left(t,\mathbf{r}^{\prime}\right)\right]_{+}=\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (104)
which is the starting point of the second quantization of the Dirac field, together with the anticommutators
$\left[\widehat{\bar{\pi}}\left(t,\mathbf{r}\right),\widehat{\pi}\left(t,\mathbf{r}^{\prime}\right)\gamma^{0}\right]_{+}=\left(\frac{\hbar c}{2}\right)^{2}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (105)
$\left[\widehat{\psi}\left(t,\mathbf{r}\right),\widehat{\pi}\left(t,\mathbf{r}^{\prime}\right)\gamma^{0}\right]_{+}=\frac{i\hbar c}{2}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (106)
$\left[\widehat{\bar{\pi}}\left(t,\mathbf{r}\right),\widehat{\psi^{{\dagger}}}\left(t,\mathbf{r}^{\prime}\right)\right]_{+}=-\frac{i\hbar c}{2}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (107)
The constraints (32) give the operator identities
$\widehat{\pi}=\frac{i\hbar c}{2}\widehat{\bar{\psi}}~{},\quad\widehat{\bar{\pi}}=-\frac{i\hbar c}{2}\widehat{\psi}~{}.$ (108)
Substituting them into Eqs. (105), (106) and (107) all reproduce Eq. (104), which is the fundamental anticommutator between the operator $\widehat{\psi}$ and its adjoint $\widehat{\psi}^{{\dagger}}$ in the quantum treatment of the Dirac field BD ; HK ; IZ .
The canonical quantization needs to be defined for arbitrary scalar phase space functions $f,g$ as well. As they are linear combinations of the type $\sum_{i}\mathcal{A}_{i}\mathcal{D}_{i}$ and the Leibniz rule holds
$\left\{f,g\right\}_{D}$ $=$ $\sum_{ij}\left\{\mathcal{A}_{i}^{\left(f\right)}\mathcal{D}_{i}^{\left(f\right)},\mathcal{A}_{j}^{\left(g\right)}\mathcal{D}_{j}^{\left(g\right)}\right\}_{D}$ (109)
$=$ $\sum_{ij}\mathcal{A}_{i}^{\left(f\right)}\left\{\mathcal{D}_{i}^{\left(f\right)},\mathcal{A}_{j}^{\left(g\right)}\right\}_{D}\mathcal{D}_{j}^{\left(g\right)}-\sum_{ij}\mathcal{A}_{j}^{\left(g\right)}\left\{\mathcal{D}_{j}^{\left(g\right)},\mathcal{A}_{i}^{\left(f\right)}\right\}_{D}\mathcal{D}_{i}^{\left(f\right)}~{}.$
The factor ordered Dirac bracket terms $\left\{\mathcal{D}_{i}^{\left(f\right)},\mathcal{D}_{j}^{\left(g\right)}\right\}_{D}$ and $\left\{\mathcal{A}_{i}^{\left(f\right)},\mathcal{A}_{j}^{\left(g\right)}\right\}_{D}$ gave no contribution (due to their second and first arguments, respectively) and we have employed the antisymmetry of the Dirac bracket in the last term. Then, the canonical quantization recipe (103) can be applied to $\left\{f,g\right\}_{D}$ as well.
The canonical quantization can also be done by starting from the reduced phase space variables (60), for which we apply the prescription (103), obtaining
$[\widehat{\psi_{1}},\widehat{\pi_{1}}\gamma^{0}]_{+}=i\hbar c\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (110)
By employing the operator identities (108), we recover the fundamental anticommutator (104) once more.
V.2 Grassmannian approach with left derivatives
The canonical quantization recipe in this case is
$\left\{F,G\right\}_{D+}^{L}\mapsto-\dfrac{i}{\hbar c}[\widehat{F},\widehat{G}\gamma^{0}]_{+}~{}.$ (111)
With the operator identities
$\widehat{\bar{\pi}}_{L}=-\dfrac{i\hbar c}{2}\widehat{\psi}~{},\quad\widehat{\pi}_{L}=-\dfrac{i\hbar c}{2}\widehat{\bar{\psi}}~{}.$ (112)
arising from the constraints (86), all quantized versions of the nontrivial Dirac brackets (97) reproduce the fundamental anticommutator (104).
V.3 Grassmannian approach with right derivatives
In this case, the canonical quantization recipe becomes
$\left\{F,G\right\}_{D+}^{R}\mapsto\dfrac{i}{\hbar c}[\widehat{F},\widehat{G}\gamma^{0}]_{+}~{}.$ (113)
The constrains (87) lead to the operator identities
$\widehat{\bar{\pi}}_{R}=\dfrac{i\hbar c}{2}\widehat{\psi}~{},\quad\widehat{\pi}_{R}=\dfrac{i\hbar c}{2}\widehat{\bar{\psi}}~{}.$ (114)
With these, the fundamental anticommutator (104) is once again reproduced from all nontrivial Dirac brackets (98).
VI Concluding Remarks
We generalized a previously successful discussion of the Schrödinger field Schrod as a constrained system through the Dirac–Bergmann algorithm to the case of the Dirac field. This was motivated by the analogy of the Lagrangians describing the Dirac field with the simplest Schiff and Hermitian Henley–Thirring Lagrangians of the Scrödinger field.
In order to follow as closely as possible the treatment of the Schrödinger field, we discussed the Dirac field classically as a spinor. Then, for the Hamiltonian evolution we introduced suitably modified, factor ordered Poisson and Dirac brackets (in which Dirac adjungate spinors are always followed by spinors) and we recovered all essential ingredients, similarly to the case of the Schrödinger field. We introduced momenta, which are Dirac adjoints of the respective spinorial field variables. This ensured that no $\gamma^{0}$ factors appear in the Poisson bracket.
The Dirac–Bergmann algorithm applied to the Hermitian Lagrangian for the Dirac field gave two primary Hamiltonian constraints, Dirac adjoints of each other. Evolving (through the factor ordered Poisson bracket) the constraints generated the Dirac equation and its Dirac adjoint as consistency conditions, holding weakly. The constraints being second class implied the introduction of the factor ordered Dirac bracket, generating time evolution. The Dirac brackets of the canonical pairs contained $1/2$ factors not present in their Poisson brackets, similarly to the case of the Schrödinger field, an essential feature in its canonical quantization.
We identified the canonical chart in which one canonical pair spans the reduced phase space while the other is the pair of constraints. In this chart the Dirac bracket on the full phase space became the Poisson bracket on the reduced phase space, on which the Hamiltonian coincides with the Hermitian Hamiltonian evaluated on shell. This feature differs from the Schrödinger case, where the reduced Hamiltonian coincides on shell with both the Schiff and Henley–Thirring Hamiltonians. For the Dirac field the Hamiltonian obtained from the simplest Lagrangian (the analogue of the Schiff Hamiltonian) carries additional contributions. The Dirac equation and its Dirac adjoint could be recovered as canonical equations on the reduced phase space.
When the Dirac field is considered an odd Grassmann variable, it obeys the same algebraic structure classically as in the quantum version of the theory. Hence, we carried out the Dirac–Bergmann algorithm for the Grassmannian Dirac field with both left and righ derivatives acting on Grassmann valued superfunctions. These involved other types of generalized Poisson and Dirac brackets, symmetric in the canonical data, rather than antisymmetric.
Finally, we provided the recipe for the canonical second quantization of all three versions of the classical Dirac dynamics. This naturally transformed all three versions of the introduced Dirac brackets into the fundamental anticommutator characterizing the second quantized Dirac field.
Appendix A The Dirac spinor represents spin $1/2$ particles
Taking any spacetime vector $V^{a}$, one can define another vector with identical components on the space spanned by the orthonormal basis $\left\{\mathrm{I}_{2},\sigma^{\mu}\right\}$:
$v$ $=$ $V^{0}\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)+V^{1}\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)+V^{2}\left(\begin{array}[]{cc}0&-i\\ i&0\end{array}\right)+V^{3}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$ (123)
$=$ $\left(\begin{array}[]{cc}V^{0}+V^{3}&V^{1}-iV^{2}\\ V^{1}+iV^{2}&V^{0}-V^{3}\end{array}\right)~{}.$ (126)
A rotation
$\left(\begin{array}[]{c}V^{0\prime}\\ V^{1\prime}\\ V^{2\prime}\\ V^{3\prime}\end{array}\right)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\cos\varphi&\sin\varphi&0\\ 0&-\sin\varphi&\cos\varphi&0\\ 0&0&0&1\end{array}\right)\left(\begin{array}[]{c}V^{0}\\ V^{1}\\ V^{2}\\ V^{3}\end{array}\right)$
with angle $\varphi$ about the $z$-axis generates
$v^{\prime}=\left(\begin{array}[]{cc}V^{0}+V^{3}&\left(V^{1}-iV^{2}\right)e^{i\varphi}\\ \left(V^{1}+iV^{2}\right)e^{-i\varphi}&V^{0}-V^{3}\end{array}\right)~{},$ (127)
thus an $SL\left(2,\mathbb{C}\right)$ transformation on $v^{a}$
$v^{\prime}=\left(\begin{array}[]{cc}e^{i\varphi/2}&0\\ 0&e^{-i\varphi/2}\end{array}\right)v\left(\begin{array}[]{cc}e^{i\varphi/2}&0\\ 0&e^{-i\varphi/2}\end{array}\right)^{{\dagger}}~{}.$ (128)
It is simple to check that $\det v=0$ is equivalent with $V^{a}$ being null and we choose $V^{a}=\frac{1}{2}\left(1,0,0,1\right)$ in order to simplify the forthcoming results. Then,
$v=\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right)=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\left(\begin{array}[]{cc}1&0\end{array}\right)~{}.$ (129)
The first factor is a left Weyl spinor $\psi_{L}$, while the second is the transposed of the Weyl adjoint $\bar{\psi}_{R}=\varepsilon\psi_{R}^{{\dagger}}$ of a right Weyl spinor $\psi_{R}$ Steane , where
$\varepsilon=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)~{}.$ (130)
Under rotations they transform as
$\psi_{L}$ $=$ $\left(\begin{array}[]{c}1\\ 0\end{array}\right)\rightarrow\psi_{L}^{\prime}=\left(\begin{array}[]{cc}e^{i\varphi/2}&0\\ 0&e^{-i\varphi/2}\end{array}\right)\left(\begin{array}[]{c}1\\ 0\end{array}\right)=e^{i\varphi/2}\left(\begin{array}[]{c}1\\ 0\end{array}\right)~{},$ (139)
$\bar{\psi}_{R}^{T}$ $=$ $\left(\begin{array}[]{cc}1&0\end{array}\right)\rightarrow\bar{\psi}_{R}^{\prime T}=\left(\begin{array}[]{cc}1&0\end{array}\right)\left(\begin{array}[]{cc}e^{i\varphi/2}&0\\ 0&e^{-i\varphi/2}\end{array}\right)^{{\dagger}}=e^{-i\varphi/2}\left(\begin{array}[]{cc}1&0\end{array}\right)~{}.$ (145)
From the latter
$\psi_{R}=\left(\begin{array}[]{cc}0&1\end{array}\right)\rightarrow\psi_{R}^{\prime}=e^{i\varphi/2}\left(\begin{array}[]{cc}0&1\end{array}\right)~{}$ (146)
can be easily deduced.
Hence, the Dirac spinor, defined through
$\psi=\left(\begin{array}[]{c}\psi^{L}\\ \psi_{R}^{T}\end{array}\right)$ (147)
transforms as
$\psi^{\prime}=e^{i\varphi/2}\psi~{}.$ (148)
This is simple proof that the Dirac spinor represents spin 1/2 particles.
Note the similarity of this proof to the electromagnetic and gravitational cases. For the electromagnetic plane wave
[…]
(with null wave vector $k^{c}$) the polarization vector $\mathfrak{e}_{a}$ in the Lorenz gauge ensuring transversality and the cancellation of the temporal and longitudinal polarizations is $\mathfrak{e}_{a}=\left(0,\mathfrak{e}_{1},\mathfrak{e}_{2},0\right)$. Under rotations about the $z$-axis with angle $\varphi$ the polarizations $\mathfrak{e}_{\pm}=\mathfrak{e}_{1}\mp i\mathfrak{e}_{2}$ transform as $\mathfrak{e}_{\pm}^{\prime}=e^{\pm i\varphi}\mathfrak{e}_{\pm}$, implying helicity ±$\pm 1$ for the photon.
A similar reasoning stands for the gravitational waves in the weak field approach. For the perturbation $h_{ab}$ of the Minkowski metric, assumed as a plane wave
[…]
the polarization tensor $\mathfrak{e}_{ab}$ in harmonic coordinates and transverse traceless gauge is
$\mathfrak{e}_{ab}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\mathfrak{e}_{11}&\mathfrak{e}_{12}&0\\ 0&\mathfrak{e}_{12}&-\mathfrak{e}_{11}&0\\ 0&0&0&0\end{array}\right)~{}.$ (151)
Under rotations about the $z$-axis with angle $\varphi$ the polarizations $\mathfrak{e}_{\pm}=\mathfrak{e}_{11}\mp i\mathfrak{e}_{12}$ transform as $\mathfrak{e}_{\pm}^{\prime}=e^{\pm 2i\varphi}\mathfrak{e}_{\pm}$, signifying helicity ±$\pm 2$ for the gravitational wave.

### v2

@@
 The Dirac equation is recovered both as consistency condition on the full phase space and as canonical equation on the reduced phase space.
-Alternatively, considering the Dirac field as odd Grassmann variable, we present the details of the Dirac–Bergmann algorithm (with either left and righ derivatives acting on Grassmann valued superfunctions and involving a different type of generalized Poisson and Dirac brackets).
+Alternatively, considering the Dirac field as odd Grassmann variable, we present the details of the Dirac–Bergmann algorithm (with either left and right derivatives acting on Grassmann valued superfunctions and involving a different type of generalized Poisson and Dirac brackets).
 We propose a recipe for the canonical second quantization of all three versions of the generalized Dirac brackets, yielding the correct fundamental anticommutator.
 I Introduction
-Most of the physical theories of interest contain constraints, for which the transition from the Lagrangian to the Hamiltonian description requires the Dirac–Bergmann algorithm, developed initially by Dirac DiracConstr and Bergmann with his collaborators (for a historical account see Salisbury and for further details SM ; Sundermayer ; HT ).
+Most of the physical theories of interest contain constraints, for which the transition from the Lagrangian to the Hamiltonian description requires the Dirac–Bergmann algorithm, developed initially by Dirac DiracConstr and Bergmann with his collaborators (for a historical account see Salisbury and for further details SM ; Sundermeyer ; HT ).
 The full machinery of the formalism through the analysis of a finite-dimensional system has been recently presented in Ref.
 Brown .
-While the corresponding discussions of the Maxwell electrodynamics and of general relativity Wald ; ADM are well known, the Schrödinger field as a constrained system was analyzed relatively late Schrod .
+Geometric generalizations of the Dirac–Bergmann algorithm have also been developed.
+The algorithm by Gotay, Nester, and Hinds in Ref.
+GNH78 employs a global presymplectic structure, generalizing and improving on the Dirac–Bergmann algorithm, as the latter can be regarded as its local version.
+In the companion paper GN79 , Gotay and Nester further emphasize that while the Dirac–Bergmann results cannot always be transferred to the velocity phase space, their method based on the presymplectic structure allows that.
+This was later investigated for a broad class of admissible Lagrangians, defined as those which allow for second-order Euler–Lagrange equations on certain submanifolds of the velocity phase space GN80 .
+A multisymplectic approach was also developed in Ref.
+multisymp and explored in the discussion of various constrained systems, including Einstein–Hilbert multisympEH and metric-affine multisympEP gravity theories, also more recently for scalar fields, Chern-Simons gravity, and bosonic string theory multisympField .
+While the Dirac–Bergmann algorithm applied for Maxwell electrodynamics or general relativity Wald ; ADM are well known, the Schrödinger field as a constrained system was analysed relatively late Schrod .
 There, the Schrödinger equation emerged both as a consistency condition in the full phase space and as canonical equation in the reduced phase space.
@@
 The Hamiltonian on the reduced phase space reproduces the Hermitian Hamiltonian evaluated on shell.
-For comparing our approach with the interpretation of the classical Dirac field as an odd Grassmann variable, in Section 4 we also carry out the Dirac–Bergmann procedure with both left and righ derivatives acting on Grassmann valued superfunctions.
+For comparing our approach with the interpretation of the classical Dirac field as an odd Grassmann variable, in Section 4 we also carry out the Dirac–Bergmann procedure with both left and right derivatives acting on Grassmann valued superfunctions.
 These involve another type of generalized Poisson and Dirac brackets, which are symmetric, rather than antisymmetric, together with definitions of the momenta different from the literature.
@@
 The suitably defined quantization recipes transform all versions of Dirac brackets in a natural way into the fundamental anticommutator of the second quantized Dirac field.
+Section 6 contains a brief comparison of our results with the flat limit of previous discussions of fermionic (Dirac) systems in gravitational field.
+Then we present our concluding remarks in Section 7.
 […]
@@
 […]
-$\mathfrak{L}_{IZ}=\frac{i\hbar c}{2}\left(\bar{\psi}\not{\partial}\psi-\not{\partial}\bar{\psi}\psi\right)-mc^{2}\bar{\psi}\psi$ (26)
+$\mathfrak{L}_{IZ}=\frac{i\hbar c}{2}\left(\bar{\psi}\not{\partial}\psi-\not{\partial}\bar{\psi}\psi\right)-mc^{2}\bar{\psi}\psi~{}.$ (26)
 Variation with respect to the pair of independent spinor variables $\bar{\psi}$ and $\psi$ give the Dirac equation (17) and its Dirac adjoint (18), respectively. The canonical momenta emerge as
@@
 III The Dirac–Bergmann algorithm
-Similarly to the Schrödinger field, either of the Lagrangians (16) or (26) characterizing the dynamics of the Dirac spinor contain linearly the generalized velocities ∂̸$\not{\partial}_{0}\psi∂$ and ∂̸$\not{\partial}_{0}\bar{\psi}∂$, a simptom of a contrained system.
+Similarly to the Schrödinger field, either of the Lagrangians (16) or (26) characterizing the dynamics of the Dirac spinor contain linearly the generalized velocities ∂̸$\not{\partial}_{0}\psi∂$ and ∂̸$\not{\partial}_{0}\bar{\psi}∂$, a simptom of a constrained system.
 Hence the Dirac–Bergmann algorithm needs to be applied for an unambiguous transition from the Lagrangian to the Hamiltonian description.
@@
 $\mathcal{H}_{C}=\frac{i\hbar c}{2}\sum_{\mu=1}^{3}\left[\left(\not{\partial}_{\mu}\bar{\psi}\right)\psi-\bar{\psi}\left(\not{\partial}_{\mu}\psi\right)\right]+mc^{2}\bar{\psi}\psi~{}$ (35)
-with a combination of the primary contraints, in which the spinor ∂̸$\not{\partial}_{0}\psi∂$ and the Dirac adjoint spinor ∂̸$\not{\partial}_{0}\bar{\psi}∂$ are unknown functions of the phase space variables.
+with a combination of the primary constraints, in which the spinor ∂̸$\not{\partial}_{0}\psi∂$ and the Dirac adjoint spinor ∂̸$\not{\partial}_{0}\bar{\psi}∂$ are unknown functions of the phase space variables.
 III.2 Factor ordered Poisson bracket
-We introduce a factor ordered Poisson bracket (defined by the requirement that in each term a Dirac adjungate spinor multiplies a spinor) of two phase space functions $f~{}$and $g$ (thus depending only of inner products of Dirac adjungate spinors with spinors) as
+We introduce a factor ordered Poisson bracket (defined by the requirement that in each term a Dirac adjoint spinor multiplies a spinor) of two phase space functions $f~{}$and $g$ (thus depending only of inner products of Dirac adjoint spinors with spinors) as
 $\left\{f\left(t,\mathbf{r}\right),g\left(t,\mathbf{r}^{\prime}\right)\right\}$ $=$ $\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi\left(t,\mathbf{r}^{\prime\prime}\right)}\right.$ (36)
@@
 The Dirac brackets (50) of the canonical phase space variables emerge as
-$\left\{\psi\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=-\frac{i}{\hbar c}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (52)
-$\left\{\pi\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{i\hbar c}{4}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{}.$ (53)
-$\left\{\psi\left(t,\mathbf{r}\right),\psi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\psi}\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{}.$ (54)
+$\left\{\psi\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=-\frac{i}{\hbar c}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (52)
+$\left\{\pi\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\frac{i\hbar c}{4}\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)~{},$ (53)
+$\left\{\psi\left(t,\mathbf{r}\right),\psi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\psi}\left(t,\mathbf{r}\right),\bar{\psi}\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{},$ (54)
 $\left\{\pi\left(t,\mathbf{r}\right),\pi\left(t,\mathbf{r}^{\prime}\right)\right\}_{D}=\left\{\bar{\pi}\left(t,\mathbf{r}\right),\bar{\pi}\left(t,\mathbf{r}^{\prime}\right)\right\}=0~{},$ (55)
@@
 $\frac{1}{2}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{1}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)$
-$-\frac{1}{2}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)$ (62)
+$-\frac{1}{2}\int\mathrm{d}\mathbf{r}^{\prime\prime}\left(\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}-\frac{\delta g\left(t,\mathbf{r}^{\prime}\right)}{\delta\psi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\frac{\delta f\left(t,\mathbf{r}\right)}{\delta\pi_{2}\left(t,\mathbf{r}^{\prime\prime}\right)}\right)~{}.$ (62)
 Thus the Dirac bracket becomes
@@
 In an alternative approach explored in Refs.
-HT and Grassmann2 , classically the Dirac field is regarded as an odd Grassmann variable.
-For comparison we explicitly present the Dirac–Bergmann algorithm for the Dirac field, treated as an odd Grassmann variable, in two equivalent formulations.
+HT and Grassmann2 , classically fermionic fields are regarded as odd Grassmann variables, nevertheless these references do not present the case of the Dirac field explicitly.
+For comparing with our earlier treatment, we explicitly present the Dirac–Bergmann algorithm for the Dirac field, treated as an odd Grassmann variable, in two equivalent formulations.
 For simplifying the notation, we also drop the arguments of the variables in this section.
@@
 $\xi^{i}\xi^{j}=-\xi^{j}\xi^{i}~{}.$ (70)
-In the above notation we have supressed the exterior or wedge product between them.
+In the above notation we have suppressed the exterior or wedge product between them.
 Such products applied repeatedly on the generators lead to superfunctions.
@@
 $\bar{\pi}_{L}$ $=$ $\dfrac{\partial^{L}\mathfrak{L}_{G}}{\partial\not{\partial}_{0}\bar{\psi}}=-\dfrac{i\hbar c}{2}\psi~{}.$ (75)
-They differ by $\gamma^{0}$ factors in the denominator from the definitions advanced in Ref.
+They differ by $\gamma^{0}$ factors in the denominator from the expressions obtainable by applying the definitions advanced in Ref.
 HT (which contain $\partial_{0}$ instead ∂̸$\not{\partial}_{0}∂$).
@@
 ∂̸$\not{\partial}_{0}\pi_{L}∂$ $=$ $-\dfrac{\partial^{L}\mathcal{H}_{L}}{\partial\psi}~{},\quad\not{\partial}_{0}\bar{\pi}_{L}=-\dfrac{\partial^{L}\mathcal{H}_{L}}{\partial\bar{\psi}}~{}.$ (77)
-After accounting for the change in the definitions of momenta, these agree with the corresponding equations presented in Ref.
+After accounting for the change in the definitions of momenta, these agree with the expressions obtainable through the canonical equations given in Ref.
 HT .
@@
 allowing to define the Dirac brackets through
-$\{F,G\}_{D+}^{L,R}=\{F,G\}_{+}^{L,R}-\int\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\{f,\phi^{i}\}_{+}^{L,R}(A_{ij}^{L,R})^{-1}\{\phi^{j},g\}_{+}^{L,R}~{}.$ (94)
+$\{F,G\}_{D+}^{L,R}=\{F,G\}_{+}^{L,R}-\int\int\mathrm{d}\mathbf{r}^{\prime\prime}\mathrm{d}\mathbf{r}^{\prime\prime\prime}\{f,\phi^{i}\}_{+}^{L,R}(A_{ij}^{L,R})^{-1}\{\phi^{j},g\}_{+}^{L,R}~{}.$ (94)
 As a result, the following two Dirac brackets emerge
@@
 $\{\psi,\pi_{R}\}_{D+}^{R}$ $=$ $\{\pi_{R},\psi\}_{D+}^{R}=\{\bar{\psi},\bar{\pi}_{R}\}_{D+}^{R}=\{\bar{\pi}_{R},\bar{\psi}\}_{D+}^{R}=-\dfrac{1}{2}\delta({\boldsymbol{r}}-{\boldsymbol{r}}^{\prime})\ .$ (98)
+The relations 97 containing left derivatives have also been given in Ref.
+Sundermeyer , however in different notation (our momenta differ by factors of $\gamma^{0}$).
 Note that now only the Dirac brackets of the canonical pairs are identical when calculated in terms of left or right derivatives.
@@
 $\left(\begin{array}[]{c}\psi_{1}^{R}=\frac{1}{2}\psi-\frac{i}{\hbar c}\bar{\pi}_{R}\\ \pi_{1}^{R}=\frac{i\hbar c}{2}\bar{\psi}+\pi_{R}\end{array}\right)$ (101)
-on the reduced phase space and the pair of contraints
+on the reduced phase space and the pair of constraints
 $\left(\begin{array}[]{c}\psi_{2}^{R}=\frac{1}{2}\psi+\frac{i}{\hbar c}\bar{\pi}_{R}=\dfrac{i}{\hbar c}\phi^{R}\\ \pi_{2}^{R}=\pi_{R}-\frac{i\hbar c}{2}\bar{\psi}=\bar{\phi}^{R}\end{array}\right)$ (102)
@@
 $\left\{F,G\right\}_{D+}^{L}\mapsto-\dfrac{i}{\hbar c}[\widehat{F},\widehat{G}\gamma^{0}]_{+}~{}.$ (111)
+After the changes in notation, this quantization scheme is equivalent to the one proposed by Casalbuoni Casalbuoni .
 With the operator identities
@@
 With these, the fundamental anticommutator (104) is once again reproduced from all nontrivial Dirac brackets (98).
-VI Concluding Remarks
+VI On the coupled gravity-Dirac field system
+Instead of a metric description, suitable to deal with bosonic (tensorial) fields, gravity can also be described in terms of 16 tetrad variables, as worked out in 1929 by Weyl Weyl29 .
+The four tetrad vectors (with 4 spacetime components each) represent local orthonormal Lorentz frames at each space-time point with respect to which spinors can be defined.
+This choice of gravitational variables naturally allows for covariant derivatives of spinors, in which the spin connection (or Ricci rotation coefficients, with 24 components, having two tetrad indices) couples to the $\gamma^{a}$ matrices.
+The canonical (Arnowitt—Deser—Misner type) formulation of the dynamics of gravity coupled to the Dirac field has been studied long time ago.
+As early as 1963 Kibble discussed the canonical form of the Lagrangian describing the interacting system of gravity and Dirac fields, in the framework of the tetrad formalism of gravity Kibble63 .
+He derived the dynamics of the Dirac field from a sum of the simplest Lagrangian density $\mathfrak{L}_{BD}$ with a spin–spin interaction Lagrangian (bearing similarity to the one arising in the Fermi interaction).
+Our analysis, by contrast, proceeds from the Hermitian Lagrangian density $\mathfrak{L}_{IZ}$.
+Kibble assumed anticommuting Dirac spinors, similarly as in our discussion employing anticommuting Grassmann variables.
+In the first part of our treatment however, we introduced factor ordering instead (requiring Dirac adjoint spinors to multiply Dirac spinors, in this order).
+Deser and Isham in Ref.
+DeserIsham76 provided a crystal-clear canonical description of the gravitational sector in terms of tetrad variables, in which the role of the six pairs of geometrodynamical variables (the induced metric components and associated momenta, the trace-modified extrinsic curvature) is taken by the twelve pairs of variables related to the spatial tetrad components and their respective associated momenta.
+While the six pairs of geometrodynamical variables can be reduced to two pairs (as the first class Hamiltonian and diffeomorphism constraints each reduce the degrees of freedom by two Sundermeyer ), in the tetrad formalism additional degrees of freedom emerge due to the six-parameter Lorentz transformations of the tetrad.
+Each of the related additional six angular momentum type constraints, being also first class, reduce the degrees of freedom by two, reestablishing two pairs of canonical variables as physical.
+The Lorentz invariance of the gravitational action was re-established by inserting the angular momentum type constraints into the action with a set of six Lagrange multipliers.
+Although the canonical Dirac action for a coupled Dirac field has not been given explicitly, no other obstacle for that was envisioned than the tedious nature of the involved calculations.
+Both the Einstein and Dirac Lagrangians, also the respective equations of motion were given in terms of a suitably chosen spinor Lagrangian in tetrads suited to the 3+1 formalism of general relativity GeheniauHenneaux77 , reproducing and generalizing earlier results by Dirac DiracGS .
+There, the spinor Lagrangian is the generalization of $\mathfrak{L}_{IZ}$ to curved spacetime, however the Hamiltonian given in the Note Added in Proof does not reproduce our result.
+In Ref.
+NelsonTeitelboim78 Nelson and Teitelboim gave the action for interacting gravitational and Dirac fields expressed in canonical form, in tetrad and spinorial variables.
+The results of Ref.
+DeserIsham76 were reproduced, supplementing the Hamiltonian and diffeomorphism constraints with the contributions from the Dirac spinor.
+The emerging Dirac Hamiltonian in the flat limit $L^{i}\mapsto\gamma^{i}$ does not agree with our Hamiltonian (29).
+Agreement can be reached only at the price of inserting the definitions of the momenta in the Hamiltonian (a forbidden step before variation), and by dropping a spurious factor $i$ from the mass term of the Hamiltonian given in Ref.
+NelsonTeitelboim78 .
+The Dirac action in Ref.
+NelsonTeitelboim78 is again the covariant version of $\mathfrak{L}_{IZ}$ (after omitting a spurious factor $i$ from the mass term), however the authors do not allow the variation with respect to the Dirac adjoint spinor.
+The discussion then proceeded in terms of the real and imaginary parts of the spinor field, hence the comparison with our methods is not immediate.
+Solely the brackets of the spinor field and its Dirac adjoint are given, reproducing our Dirac bracket (52) in the flat limit, however the Dirac brackets (57-58) between the canonical variables, which are the most important in our discussion, are missing.
+Therefore, our discussion goes beyond the flat spacetime limit of the results of the references mentioned above.
+It is well known that gravity can be also described in terms of torsion Hayashi ; Hayashi2 or nonmetricity NewerGR , recently summarized in Ref.
+Heinsenberg1 .
+The three approaches are sometimes referred as the geometric trinity of gravity Trinity .
+It has been proven by Koivisto, that the Dirac fermions described by the covariant version of the Hermitian Lagrangian $\mathcal{L_{IZ}}$ are unaffected by non-metricity Koivisto .
+Moreover, the consistency of the coupling to the torsion has been disputed Obukhov .
+Hence, when the connection does not have a contorsion part, but it has a disformation part (related to its nonmetricity $Q_{abc}$), Dirac fermions do not couple to the latter and their minimal coupling to any torsionfree gravity theory is viable HeisenbergJCAP ; telecoupling .
+Modified gravity generalizations of both $f(R)$ (see the monography HarkoLobo with references therein, also LeviSaid ) and $f(Q)$ types (see Refs.
+LazkozQ ; BoehmerQ ; HeisenbergQ ) have been discussed extensively.
+Dirac fields in $f(Q)$ theories were investigated in Ref.
+CarloniQ .
+In principle, $f(R,Q)$ type gravity models could also properly incorporate minimally coupled Dirac spinors.
+VII Concluding Remarks
 We generalized a previously successful discussion of the Schrödinger field Schrod as a constrained system through the Dirac–Bergmann algorithm to the case of the Dirac field.
-This was motivated by the analogy of the Lagrangians describing the Dirac field with the simplest Schiff and Hermitian Henley–Thirring Lagrangians of the Scrödinger field.
+This was motivated by the analogy of the Lagrangians describing the Dirac field with the simplest Schiff and Hermitian Henley–Thirring Lagrangians of the Schrödinger field.
 In order to follow as closely as possible the treatment of the Schrödinger field, we discussed the Dirac field classically as a spinor.
-Then, for the Hamiltonian evolution we introduced suitably modified, factor ordered Poisson and Dirac brackets (in which Dirac adjungate spinors are always followed by spinors) and we recovered all essential ingredients, similarly to the case of the Schrödinger field.
+Then, for the Hamiltonian evolution we introduced suitably modified, factor ordered Poisson and Dirac brackets (in which Dirac adjoint spinors are always followed by spinors) and we recovered all essential ingredients, similarly to the case of the Schrödinger field.
 We introduced momenta, which are Dirac adjoints of the respective spinorial field variables.
@@
 When the Dirac field is considered an odd Grassmann variable, it obeys the same algebraic structure classically as in the quantum version of the theory.
-Hence, we carried out the Dirac–Bergmann algorithm for the Grassmannian Dirac field with both left and righ derivatives acting on Grassmann valued superfunctions.
+Hence, we carried out the Dirac–Bergmann algorithm for the Grassmannian Dirac field with both left and right derivatives acting on Grassmann valued superfunctions.
 These involved other types of generalized Poisson and Dirac brackets, symmetric in the canonical data, rather than antisymmetric.